OpenStudy (anonymous):

Find the exact value of tan 255º.

OpenStudy (jdoe0001):

keep in mind that 255 = 135 + 120 so \(\bf tan(\alpha+\beta) = \cfrac{tan(\alpha)+tan(\beta)}{1-tan(\alpha)tan(\beta)}\\ \quad \\ tan(255^o)\implies tan(135^o+120^o)=\cfrac{tan(135^o)+tan(120^o)}{1-tan(135^o)tan(120^o)}\) and you can get those angles from your Unit Circle

OpenStudy (anonymous):

okay that looks familiar. but what would be my next step, canceling things out or dividing?

OpenStudy (jdoe0001):

no.... just get the tangent values, they'd be fractions or just an a radical and add them up and simplify the rational since the tangents will give you a number, you'd end up with a number, either fractional or integer

OpenStudy (anonymous):

would it look like this \[3\pi/4+2\pi/3 \over 1- 3\pi/4 * 2\pi/3\]

OpenStudy (anonymous):

scratch what i just said, it's wrong. i know tangent is y/x and i found out that tangent of 120 degrees is - square root of 3, but how would you get that?

OpenStudy (jdoe0001):

hmmm how do we get tangent of 120 degrees is \(\bf -\sqrt{3}\) ?

OpenStudy (jdoe0001):

say for example... what's the tangent now of 135 degrees?

OpenStudy (anonymous):

would it be 1 or -1?

OpenStudy (jdoe0001):

same/same = 1 -same/same = -1 same/-same = -1

OpenStudy (anonymous):

so -1

OpenStudy (anonymous):

i think i see how it's - square root of 3 the 2s cancel out and so does the one, so its - square root of 3?

OpenStudy (jdoe0001):

so is -1... ok, now we know the tangents of 135 and 120 degrees, we know that thus \(\bf tan(135^o+120^o)=\cfrac{tan(135^o)+tan(120^o)}{1-tan(135^o)tan(120^o)}\\ \quad \\ \implies tan(135^o+120^o)=\cfrac{(-1)+(-\sqrt{3})}{1-(-1)(-\sqrt{3})}\)

OpenStudy (anonymous):

okay i get that. but i was never good at simplifying. would the numerator be square root of 3?

OpenStudy (jdoe0001):

no, it'll be just a simple addition nothing to cancel out as far as I can tell you may just be asked to leave it rational too, and that'd be ok

OpenStudy (anonymous):

i think i got it. would it be \[2 + \sqrt{3}\]

OpenStudy (jdoe0001):

\(\bf \cfrac{(-1)+(-\sqrt{3})}{1-(-1)(-\sqrt{3})}\implies \cfrac{-1-\sqrt{3}}{1-\sqrt{3}}\)

OpenStudy (jdoe0001):

can't add the radical to the integer, unless you make it a real-number,